System Of Two Linear Equation Question : A bookstore sold 18 books for a total of $176 . If some of the books were sold for $7 and some for $12 how many of each are there?
Answer by Cromlix(4381)

Hi there,
Make $7 books = x
and $12 books = y
x + y = 18.......(1)
7x + 12y = 176...(2)
Multiply (1) by 12
12x + 12y = 216....(1)
7x + 12y = 176.....(2)
Take Eq(2) from Eq(1)
5x = 40
x = 8
Substitute x = 8 into Eq(1)
x + y = 18
8 + y = 18
y = 18 - 8
y = 10
8 x $7 and 10 x $12 books were sold.
Hope this helps :-)